(2x^(2)+5x-3)/(x^(2)+x-6)÷(3x^(2)+4x+1)/(3x^(2)-5x-2) simplified
Answer by DrBeeee(684)

Factor all quadratics
(1) 2x^2 +5x -3 = (2x-1)(x+3)
(2) x^2 +x -6 = (x+3)(x-2)
(3) 3x^2 +4x +1 = (3x+1)(x+1)
(4) 3x^2 -5x -2 = (3x+1)(x-2)
We are asked to simplify
(5) ((1)/(2))/((3)/(4)) which we flip and multiply
(6) ((1)/(2))*((4)/(3)) or
(7) ((2x-1)/(x-2))*((x-2)/(x+1)) which reduces to
(8) (2x-1)/(x+1); x not = {-3,-1,-1/3,2}